mole of electron present in 51g of nh3 is 
1)9
2)21
3)3
4)30

Grade:12th pass

2 Answers

Mahima Kanawat
5 years ago
Dear student
Moles of NH3  = 51/17 =3
No of electrons in 1 mole NH3 = 7 in N + 3*1 in H =10
Therefore, total moles of electrons in 3 mole NH3= 3*10 = 30
WITH REGARDS 
MAHIMA 
ASKIITIANS FORUM EXPERT 
Manoj Singh
5 years ago
As we know
n=m/M so no. of moles of NH3 is 3 mol
But,we have to calculate no of electrons we multiply 3 mol by no of electrons in NH3
No of electrons in NH3 = 7(N)+3(H3)=10 electrons 
So 1 mol= 10 electrons
      3 mol= 30 electrons
Answer: 30